# Light Dirac neutrino portal dark matter with gauged $U(1)_{B-L}$ symmetry

### 11 Dec 2023 v1

@@ IV WIMP type dark matter
If the Yukawa coupling among the dark sector particles i.e. $\phi_{1}$, $\psi$ and $\nu_{R}$ is large ($y_{\phi_{1}}>10^{-7}$), the dark sector can be in thermal equilibrium among themselves even after their decoupling from the SM bath. Since the thermalisation of $\phi_{1},\nu_{R}$ with the SM bath relies on Higgs portal and $U(1)_{B-L}$ portal couplings instead of Yukawa $y_{\phi_{1}}$, we have two different sub-cases depending upon whether the Yukawa interactions in dark sector go out of equilibrium before or after the dark sector decouples from the thermal bath. Among the dark sector particles, $\phi_{1}$ and $\nu_{R}$ can have interactions with other particles in the bath bath via the processes - $\phi_{1}X\to\phi_{1}X$, $\phi_{1}Z^{\prime}\to\phi_{1}Z^{\prime}$ and $\nu_{R}X\to\nu_{R}X$, here $X$ denotes SM particles. Since $Z^{\prime}$ has sizeable interactions with all the SM fermions, we consider it to be in the bath while dark sector particles decouple. While the first two processes can arise via contact interactions too, the $\nu_{R}$ scattering arises via mediation of heavy $Z^{\prime}$ boson only. Unlike other interactions, $\nu_{R}$ interactions with the bath rely significantly on resonant enhancement and hence we consider $\nu_{R}\bar{\nu}_{R}\rightarrow X\bar{X}$ via resonantly enhanced s-channel mediation of $Z^{\prime}$ to calculate its decoupling Abazajian:2019oqj . Therefore, the ratio of interaction rate of dark sectors with SM bath to the expansion rate, $\mathcal{H}$, can be written as Gondolo:2012vh ; Heeck:2014zfa
@@
and $\xi$ is the ratio of dark sector temperature to that of the SM bath, $\frac{T_{\nu_{R}}}{T}$ and $\alpha=6\times\frac{7}{8}\frac{\pi^{2}}{30}$. After decoupling, $\langle\sigma v\rangle_{\rm eff}$ in Eq. (18) can depend upon both $x$ as well as $\xi$. We solve Eq. (18) and Eq. (20) simultaneously from $T=T_{\rm dec}$ to the onset of the BBN era namely, $T\sim 10$ MeV and estimate the corresponding DM relic and $\Delta N_{\rm eff}$. Fig. 8 shows the evolution of $T_{\nu_{R}}$ for different benchmark parameters. Different $T_{\nu_{R}}$ can lead to changes in $\Delta N_{\rm eff}$. The first term on the RHS of Eq. (20) is a positive term as $Y\geq Y^{\rm eq}$ for $T<T_{\rm dec}$. Thus, this term indicates the increase in $T_{\nu_{R}}$ over $T$ due to annihilation of heavy dark sector particles $\phi_{1}$ and $\psi$ into $\nu_{R}$. On the other hand, the second term on the RHS of the same equation indicates the decrease in $T_{\nu_{R}}$ over $T$ due to changes in entropy degrees of freedom $g^{s}_{*}$. The combined effects of these two terms can be seen in Fig. 8. For an analytical description of the behaviour of the plots shown in Fig. 8, please refer to appendix B.
@@
The left panel plot in Fig. 8 shows the variation of $\xi^{4}$ with respect to bath temperature $T$ for different choices of $m_{Z^{\prime}}$. The rest of the parameter are kept fixed as shown in the plot. Increasing $m_{Z^{\prime}}$ gives a higher decoupling temperature, $T_{\rm dec}$. As long as $T_{\rm dec}\lesssim m_{\phi_{1}}$, a higher decoupling temperature leads to a larger conversion of $\phi_{1}$ and $\psi$ to $\nu_{R}$ resulting in a larger $\frac{T_{\nu_{R}}}{T}$. Another way to understand it is from entropy conservation (see appendix B). For a higher $T_{\rm dec}$, we have a larger dark sector entropy degrees of freedom $g^{\rm s,DS}_{*}$ resulting in a larger $\frac{T_{\nu_{R}}}{T}$. From an analytical estimation, we get a maximum value of $\xi^{4}$ to be 2.6 provided $g^{s}_{*}$ remains same from the epoch of dark sector decoupling to DM freeze-out, as shown in appendix B. This is consistent with the numerical results shown in left panel of Fig. 8. The reason why maximum value of $\xi^{4}$ has not reached $2.6$ in the figure is that $g^{s}_{*}$ changes from the period of kinetic decoupling of dark sector to DM freeze-out. The middle panel plot shows the results for different values of dark sector Yukawa coupling. As the dark sector Yukawa coupling does not affect the DM-SM decoupling temperature, so all the coloured lines correspond to the same decoupling temperature. For a larger Yukawa coupling, conversion of $\phi_{1}$ and $\psi$ to $\nu_{R}$ is possible and as a result, temperature of $\nu_{R}$ with respect to SM bath increases. For the green line in the middle panel plot of Fig. 8, the dark sector freeze-out occurs before $T_{\rm dec}$ due to small Yukawa coupling. Hence the first term on the RHS of Eq. (20) ceases and due to the second term, the ratio $\frac{T_{\nu_{R}}}{T}$ decreases. To see the effect of $m_{\psi}$, in the right panel plot, we vary $m_{\psi}$ keeping the other parameter constant. For a smaller $m_{\psi}$, freeze-out occurs late, hence we get a higher $\frac{T_{\nu_{R}}}{T}$ ratio.
@@
In the plots shown in Fig. 10, we show the parameter space in $m_{Z^{\prime}}$ versus $g_{BL}$ plane for WIMP type DM. Each of these plots correspond to a fixed value of $\lambda_{H\phi_{1}}$. The other parameters fixed in all the plots are, $m_{\phi_{1}}=60$ GeV, $m_{\psi}=15$ GeV, $y_{\phi_{1}}=0.28$. The solid black line separates regime 1 (upper left triangular region) from regime 2. The solid brown line separates the regime where kinetic decoupling temperature is determined by Higgs portal coupling from the one where it is determined by the gauge portal coupling. In the grey coloured region below the solid brown line, $g_{BL}$ and $m_{Z^{\prime}}$ have no role in the kinetic decoupling of DM. In the upper panel plots, the region below the magenta line is where we have $\Delta N_{\rm eff}>0.28$, disfavoured by Planck 2018 limits. In the grey coloured region, due to constant decoupling temperature, we have constant $\Delta N_{\rm eff}$. Due to smaller Higgs-portal coupling in the upper right panel plot compared to the upper left one, the grey coloured region shrinks. In the lower left panel plot, we increase the Higgs-portal coupling compared to upper left panel plot to $2.5\times 10^{-4}$. For this particular choice of $\lambda_{H\phi_{1}}$, the region between the black and brown line has $\Delta N_{\rm eff}<0.28$ and the region below the brown line has $\Delta N_{\rm eff}$=0.28, which is the maximum allowed value from Planck 2018 limits. In the lower right panel plot, $\lambda_{H\phi_{1}}$ is increased even further to $10^{-3}$ giving $\Delta N_{\rm eff}<0.28$ in the whole regime 1. For regime 2, as DM freezes out before RHN decoupling, the bound from $\Delta N_{\rm eff}$ (denoted by region above the green line) is same as the bound shown in Fig. 1. The entire parameter space shown in these plots satisfy the criteria of DM relic abundance. This is because DM abundance has very weak dependence on the kinetic decoupling temperature. For the WIMP type DM scenario, we also check the constraints from direct detection and find them to be very weak due to radiative suppression. The details are given in appendix C.
@@ Acknowledgements
Appendix A Calculation of FSL
The free-streaming length (FSL) can be quantified as 2012MNRAS.424..684S-
@@ Appendix A Calculation of FSL
Appendix B Estimation of dark sector temperature
@@
However, we can have a value of $g^{\rm s,DS}_{\rm*,dec}<10.75$ at the epoch of decoupling. This is because before decoupling, some fraction of dark sector particles can become non-relativistic. For example, let us consider $m_{\phi_{1}}=60$ GeV, $m_{\psi}=15$ GeV. Decoupling above $T\sim 60$ GeV gives $g^{s,\rm DS}_{*,\rm dec}<10.75$ whereas decoupling after $T\sim 60$ GeV (e.g. $10$ GeV) gives $g^{\rm s,DS}_{*,\rm dec}<10.75$. This results in a smaller increase in dark sector temperature from SM bath. The same behaviour can also been seen from Fig. 8.
Appendix C Direct Detection of WIMP type DM

### 09 Apr 2024 v2

@@ IV WIMP type dark matter
If the Yukawa coupling among the dark sector particles i.e. $\phi_{1}$, $\psi$ and $\nu_{R}$ is sufficiently strong ($y_{\phi_{1}}\gg 10^{-7}$), the dark sector can be in thermal equilibrium among themselves even after their decoupling from the SM bath. Since the thermalisation of $\phi_{1},\nu_{R}$ with the SM bath relies on Higgs portal and $U(1)_{B-L}$ portal couplings instead of Yukawa $y_{\phi_{1}}$, we have two different sub-cases depending upon whether the Yukawa interactions in dark sector go out of equilibrium before or after the dark sector decouples from the thermal bath. Among the dark sector particles, $\phi_{1}$ and $\nu_{R}$ can have interactions with other particles in the bath via the processes - $\phi_{1}X\to\phi_{1}X$, $\phi_{1}Z^{\prime}\to\phi_{1}Z^{\prime}$ and $\nu_{R}X\to\nu_{R}X$, here $X$ denotes SM particles. Since $Z^{\prime}$ has sizeable interactions with all the SM fermions, we consider it to be in the bath while dark sector particles decouple. While the first two processes can arise via contact interactions too, the $\nu_{R}$ scattering arises via mediation of heavy $Z^{\prime}$ boson only. Unlike other interactions, $\nu_{R}$ interactions with the bath rely significantly on resonant enhancement and hence we consider $\nu_{R}\bar{\nu}_{R}\rightarrow X\bar{X}$ via resonantly enhanced s-channel mediation of $Z^{\prime}$ to calculate its decoupling Abazajian:2019oqj . Therefore, the ratio of interaction rate of dark sectors with SM bath to the expansion rate, $\mathcal{H}$, can be written as Gondolo:2012vh ; Heeck:2014zfa
@@
and $\xi$ is the ratio of dark sector temperature to that of the SM bath, $\frac{T_{\nu_{R}}}{T}$ and $\alpha=6\times\frac{7}{8}\frac{\pi^{2}}{30}$. After decoupling, $\langle\sigma v\rangle_{\rm eff}$ in Eq. (19) can depend upon both $x$ as well as $\xi$. We solve Eq. (19) and Eq. (21) simultaneously from $T=T_{\rm dec}$ to the onset of the BBN era namely, $T\sim 10$ MeV and estimate the corresponding DM relic and $\Delta N_{\rm eff}$. Fig. 8 shows the evolution of $T_{\nu_{R}}$ for different benchmark parameters. Different $T_{\nu_{R}}$ can lead to changes in $\Delta N_{\rm eff}$. The first term on the RHS of Eq. (21) is a positive term as $Y\geq Y^{\rm eq}$ for $T<T_{\rm dec}$. Thus, this term indicates the increase in $T_{\nu_{R}}$ over $T$ due to annihilation of heavy dark sector particles $\phi_{1}$ and $\psi$ into $\nu_{R}$. On the other hand, the second term on the RHS of the same equation indicates the decrease in $T_{\nu_{R}}$ over $T$ due to changes in entropy degrees of freedom $g^{s}_{*}$. The combined effects of these two terms can be seen in Fig. 8. For an analytical description of the behaviour of the plots shown in Fig. 8, please refer to appendix D.
@@
The left panel plot in Fig. 8 shows the variation of $\xi^{4}$ with respect to bath temperature $T$ for different choices of $m_{Z^{\prime}}$. The rest of the parameter are kept fixed as shown in the plot. Increasing $m_{Z^{\prime}}$ gives a higher decoupling temperature, $T_{\rm dec}$. As long as $T_{\rm dec}\lesssim m_{\phi_{1}}$, a higher decoupling temperature leads to a larger conversion of $\phi_{1}$ and $\psi$ to $\nu_{R}$ resulting in a larger $\frac{T_{\nu_{R}}}{T}$. Another way to understand it is from entropy conservation (see appendix D). For a higher $T_{\rm dec}$, we have a larger dark sector entropy degrees of freedom $g^{\rm s,DS}_{*}$ resulting in a larger $\frac{T_{\nu_{R}}}{T}$. From an analytical estimation, we get a maximum value of $\xi^{4}$ to be 2.6 provided $g^{s}_{*}$ remains same from the epoch of dark sector decoupling to DM freeze-out, as shown in appendix D. This is consistent with the numerical results shown in left panel of Fig. 8. The reason why maximum value of $\xi^{4}$ has not reached $2.6$ in the figure is that $g^{s}_{*}$ changes from the period of kinetic decoupling of dark sector to DM freeze-out. The middle panel plot shows the results for different values of dark sector Yukawa coupling. As the dark sector Yukawa coupling does not affect the DM-SM decoupling temperature, so all the coloured lines correspond to the same decoupling temperature. For a larger Yukawa coupling, conversion of $\phi_{1}$ and $\psi$ to $\nu_{R}$ is possible and as a result, temperature of $\nu_{R}$ with respect to SM bath increases. For the green line in the middle panel plot of Fig. 8, the dark sector freeze-out occurs before $T_{\rm dec}$ due to small Yukawa coupling. Hence the first term on the RHS of Eq. (21) ceases and due to the second term, the ratio $\frac{T_{\nu_{R}}}{T}$ decreases. To see the effect of $m_{\psi}$, in the right panel plot, we vary $m_{\psi}$ keeping the other parameter constant. For a smaller $m_{\psi}$, freeze-out occurs late, hence we get a higher $\frac{T_{\nu_{R}}}{T}$ ratio.
@@
In the plots shown in Fig. 10, we show the parameter space in $m_{Z^{\prime}}$ versus $g_{BL}$ plane for WIMP type DM. Each of these plots correspond to a fixed value of $\lambda_{H\phi_{1}}$. The other parameters fixed in all the plots are, $m_{\phi_{1}}=60$ GeV, $m_{\psi}=15$ GeV, $y_{\phi_{1}}=0.28$. The solid black line separates regime 1 (upper left triangular region) from regime 2. The solid brown line separates the regime where kinetic decoupling temperature is determined by Higgs portal coupling from the one where it is determined by the gauge portal coupling. In the grey coloured region below the solid brown line, $g_{BL}$ and $m_{Z^{\prime}}$ have no role in the kinetic decoupling of dark sector. In the upper panel plots, the region below the magenta line is where we have $\Delta N_{\rm eff}>0.28$, disfavoured by Planck 2018 limits. In the grey coloured region, due to constant decoupling temperature, we have constant $\Delta N_{\rm eff}$. Due to smaller Higgs-portal coupling in the upper right panel plot compared to the upper left one, the grey coloured region shrinks. In the lower left panel plot, we increase the Higgs-portal coupling compared to upper left panel plot to $2.5\times 10^{-4}$. For this particular choice of $\lambda_{H\phi_{1}}$, the region between the black and brown line has $\Delta N_{\rm eff}<0.28$ and the region below the brown line has $\Delta N_{\rm eff}$=0.28, which is the maximum allowed value from Planck 2018 limits. In the lower right panel plot, $\lambda_{H\phi_{1}}$ is increased even further to $10^{-3}$ giving $\Delta N_{\rm eff}<0.28$ in the whole regime 1. For regime 2, as DM freezes out before RHN decoupling, the bound from $\Delta N_{\rm eff}$ (denoted by region above the green line) is same as the bound shown in Fig. 1. The entire parameter space shown in these plots satisfy the criteria of DM relic abundance. This is because DM abundance has very weak dependence on the kinetic decoupling temperature. For the WIMP type DM scenario, we also check the constraints from direct detection and find them to be very weak due to radiative suppression. The details are given in appendix E.
@@ Acknowledgements
Appendix A Anomaly cancellation in gauged $B-L$ model
$U(1)_{B-L}$ gauge symmetry with only the SM fermions is not anomaly free. This is because the triangle anomalies for both $U(1)^{3}_{B-L}$ and the mixed $U(1)_{B-L}-(\text{gravity})^{2}$ diagrams are non vanishing. These triangle anomalies for the SM fermion content are given as
$\mathcal{A}_{1}\left[U(1)^{3}_{B-L}\right]=\mathcal{A}^{\text{SM}}_{1}\left[U(1)^{3}_{B-L}\right]=-3\,,$
$\mathcal{A}_{2}\left[(\text{gravity})^{2}\times U(1)_{B-L}\right]=\mathcal{A}^{\text{SM}}_{2}\left[(\text{gravity})^{2}\times U(1)_{B-L}\right]=-3\,.$ (24)
If three right handed neutrinos are added to the model, they contribute $\mathcal{A}^{\text{New}}_{1}\left[U(1)^{3}_{B-L}\right]=3,\mathcal{A}^{\text{New}}_{2}\left[(\text{gravity})^{2}\times U(1)_{B-L}\right]=3$ leading to vanishing total of triangle anomalies. This is the most natural and economical $U(1)_{B-L}$ model studied extensively in the literature. However, there exist non-minimal ways of constructing anomaly free versions of $U(1)_{B-L}$ model. For example, it has been known for a few years that three right handed neutrinos with $B-L$ charges $5,-4,-4$ can also give rise to vanishing triangle anomalies Montero:2007cd as follows.
$\mathcal{A}_{1}\left[U(1)^{3}_{B-L}\right]=\mathcal{A}^{\text{SM}}_{1}\left[U(1)^{3}_{B-L}\right]+\mathcal{A}^{\text{New}}_{1}\left[U(1)^{3}_{B-L}\right]=-3+\left[-5^{3}-(-4)^{3}-(-4)^{3}\right]=0\,,$
$\mathcal{A}_{2}\left[(\text{gravity})^{2}\times U(1)_{B-L}\right]$ $=\mathcal{A}^{\text{SM}}_{2}\left[(\text{gravity})^{2}\times U(1)_{B-L}\right]+\mathcal{A}^{\text{New}}_{2}\left[(\text{gravity})^{2}\times U(1)_{B-L}\right]\,,$
$=-3+\left[-5-(-4)-(-4)\right]=0\,.$ (25)
Another solution to the anomaly cancellation conditions with irrational $B-L$ charges of new fermions was proposed by the authors of Wang:2015saa where both DM and neutrino mass can have a common origin through radiative linear seesaw. Very recently, another anomaly free $U(1)_{B-L}$ framework was proposed where the additional right handed fermions possess more exotic $B-L$ charges namely, $-4/3,-1/3,-2/3,-2/3$ Patra:2016ofq . These four chiral fermions constitute two Dirac fermion mass eigenstates, the lighter of which becomes the DM candidate having either thermal Patra:2016ofq or non-thermal origins Biswas:2016iyh . The light neutrino mass in this model arises from type II seesaw mechanism with the new chiral fermions playing no role in it. In Nanda:2017bmi , such chiral fermions with fractional charges were also responsible for generating light neutrino masses at one loop level. In the recent work on $U(1)_{B-L}$ gauge symmetry with two component DM Bernal:2018aon , the authors considered two right handed neutrinos with $B-L$ number -1 each so that the model still remains anomalous. The remaining anomalies were cancelled by four chiral fermions with fractional $B-L$ charges leading to two Dirac fermion mass eigenstates both of which are stable and hence DM candidates. In Biswas:2019ygr , type III seesaw was implemented in in $U(1)_{B-L}$ model while introducing additional chiral fermions to keep the model anomaly free. Appropriate choice of chiral fermions also leads to multi-component DM in such $U(1)_{B-L}$ symmetric type III seesaw scenario.
While it is possible to study Dirac neutrino portal DM in gauged $U(1)_{B-L}$ frameworks with different anomaly free combinations of chiral fermions, it will require additional scalar fields in order to be consistent with the desired DM phenomenology and non-zero Dirac neutrino mass. For example, if we consider three right handed neutrinos with $B-L$ charges $5,-4,-4$, we will require two additional Higgs doublets of $U(1)_{B-L}$ charge $6,3$ respectively in order to generate Dirac neutrino masses at tree level. Two singlet scalars $\phi_{1,2}$ and DM $\psi$, similar to our model, can be incorporated to generate the Dirac neutrino portal of DM and break $U(1)_{B-L}$ symmetry spontaneously. Similar non-minimal scalar content will be required for other chiral fermions mentioned above. While the calculations will be more involved due to more particles, the generic conclusions reached in our work should not change significantly.
Appendix B Calculation of FSL
The free-streaming length (FSL) can be quantified as Schneider:2011yu -
@@ Appendix B Calculation of FSL
Appendix C Calculation of LHC bound on $g_{BL}-m_{Z^{\prime}}$ plane
We evaluate the constraint on $U(1)_{B-L}$ gauge coupling $g_{BL}$ from dilepton channel: $pp\to Z^{\prime}\to l^{+}l^{-}$. Using the narrow-width approximation, the cross-section $pp\to Z^{\prime}$ can be written as
[…]
where $\hat{s}=xys$ with $f_{q}$, $f_{\bar{q}}$ being the parton distribution functions for quarks and anti-quark respectively. The narrow-width approximation cross section $\hat{\sigma}(\hat{s})$ is given as
$\hat{\sigma}(\hat{s})=\frac{4\pi^{2}}{3}\frac{\Gamma(Z^{\prime}\to q\bar{q})}{m_{Z^{\prime}}}\delta(\hat{s}-m_{Z^{\prime}}^{2}).$ (36)
We use the package CalcHEP for the computation of $\sigma(pp\to Z^{\prime})$ where CTEQ6L is used for parton distribution functions.
In our model, apart from quarks and leptons, three $\nu_{R}$s and two scalars $\phi_{1}$ and $\phi_{2}$ are charged under $U(1)_{B-L}$. The decay of $Z^{\prime}$ to both the BSM scalars $\phi_{1}$ and $\phi_{2}$ is not possible due to kinematic restrictions imposed. The decay width of $Z^{\prime}$ to quarks, charged leptons, SM neutrinos and $\nu_{R}$ can be written as
$\Gamma(Z^{\prime}\to q\bar{q})\simeq 12\times 3\times\frac{1}{9}\frac{g^{2}_{BL}}{24\pi}m_{Z^{\prime}}=4\times\frac{g^{2}_{BL}}{24\pi}m_{Z^{\prime}}$ (37)
$\Gamma(Z^{\prime}\to l^{-}l^{+})\simeq 6\times\frac{g^{2}_{BL}}{24\pi}m_{Z^{\prime}}$
$\Gamma(Z^{\prime}\to\nu_{L}\bar{\nu}_{L})\simeq 3\times\frac{g^{2}_{BL}}{24\pi}m_{Z^{\prime}}$
$\Gamma(Z^{\prime}\to\nu_{R}\bar{\nu}_{R})\simeq 3\times\frac{g^{2}_{BL}}{24\pi}m_{Z^{\prime}}$
respectively, assuming massless final states. With this, we can calculate the cross-section $\sigma(pp\to Z^{\prime}\to l^{+}l^{-})\approx\sigma(pp\to Z^{\prime})\,{\rm BR}(Z^{\prime}\to l^{-}l^{+})$ for particular values of $m_{Z^{\prime}}$ and $g_{BL}$. As $\sigma(pp\to Z^{\prime})$ is proportional to $g^{2}_{BL}$, we can normalise it by dividing with $g^{2}_{BL}$.
Figure 11: Left panel: $\frac{\sigma(pp\to Z^{\prime})\,{\rm BR}(Z^{\prime}\to l^{-}l^{+})}{g^{2}_{BL}}$ vs $m_{Z^{\prime}}$. Right panel: Bound on $g_{BL}$ vs $m_{Z^{\prime}}$ plane.
The left panel plot of Fig. 11 shows $\frac{\sigma(pp\to Z^{\prime})\,{\rm BR}(Z^{\prime}\to l^{-}l^{+})}{g^{2}_{BL}}$ as a function of $m_{Z^{\prime}}$ for our model. In the right panel plot, we show the bound on $U(1)_{B-L}$ gauge coupling for our model in magenta region. The bound is obtained by comparing the left plot with the final results from ATLAS collaboration of the LHC Run-2 with $139$ $\text{fb}^{-1}$ integrated luminosity.
Appendix D Estimation of dark sector temperature
@@
However, we can have a value of $g^{\rm s,DS}_{\rm*,dec}<10.75$ at the epoch of decoupling. This is because before decoupling, some fraction of dark sector particles can become non-relativistic. For example, let us consider $m_{\phi_{1}}=60$ GeV, $m_{\psi}=15$ GeV. Decoupling above $T\sim 60$ GeV gives $g^{s,\rm DS}_{*,\rm dec}\sim 10.75$ whereas decoupling after $T\sim 60$ GeV (e.g. $10$ GeV) gives $g^{\rm s,DS}_{*,\rm dec}<10.75$. This results in a smaller increase in dark sector temperature from SM bath. The same behaviour can also been seen from Fig. 8.
Appendix E Direct Detection of WIMP type DM